OpenStudy (anonymous):

five sqre root of y minus six = 5

OpenStudy (anonymous):

Is it: \[\sqrt[5]{y-6}=5\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

how do you get that fomat of type?

OpenStudy (anonymous):

If so, then bring the \[\sqrt[5]{?}\] over to the other side of the equals sign, so you have: y-6=5^5. Then you bring the six over to the other side to make: y=(5^5)+6 So, y= 3125+6 = 3131.

OpenStudy (anonymous):

Click on equation just below where you type.

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

I get that you have to take the 5 squre root to the other side, and I get that you have to bring the six over, but I'm confused after that.

OpenStudy (anonymous):

How are you confused? That's all you need to do.

OpenStudy (anonymous):

Do you put a negative in front of the square root because you are switching sides?

OpenStudy (anonymous):

I also don't understand how the y= 3125 + 6 = 3131

OpenStudy (anonymous):

No, the square root sign changes to a square sign. For example, \[\sqrt[5]{x}\] changes to \[x^{5}\]

OpenStudy (anonymous):

Because 5x5x5x5x5=3125, then you add the six on to it to get y.

OpenStudy (anonymous):

So when you change sides with a square root ( y square root x) you take x and times it by the power of y?

OpenStudy (anonymous):

Basically, If you had \[\sqrt[5]{6+y}=3\] then the \[\sqrt[5]{6+y}\] would change to a \[3^{2}\] when it moves from the left hand side of the equals sign to the right hand side. Do you get it now?

OpenStudy (anonymous):

Sorry, I mean \[3^{5} \]

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

XD Basically, what you said was correct.

OpenStudy (anonymous):

Wait, but if you have 5 squre root of 6+y = 3, then you would do 6 to the power of 5 then do y to the power of 5?

OpenStudy (anonymous):

No, because your moving the power of 5 to the other side of the equation. So you'll have a 3 to the power of 5.